OpenStudy (anonymous):

Calculate the average speed for each of the following. 126 km in 2 1/4 h. ________________ 292 1/2 km in 4 1/2 h. ________________ 288 km in 3 3/5 h. _____________________ Can someone provide the answer and explain how you got the answer so I understand?

OpenStudy (anonymous):

your job in each of these is to divide

OpenStudy (anonymous):

so for the first one you need \[126\div 2\tfrac{1}{4}\]

OpenStudy (anonymous):

and if you want to divide, you need to take that stupid mixed number and turn it in to an improper faction. so \[2\tfrac{1}{4}=\frac{2\times 4+1}{4}=\frac{9}{4}\]

OpenStudy (anonymous):

you divide by a fraction by multiplying by the reciprocal, so the first problem is solve by writing \[126\div \frac{9}{4}=126\times \frac{4}{9}=14\times 4=56\]

OpenStudy (anonymous):

second two are similar. try them and if you can't get them ask

OpenStudy (anonymous):

Thanks a lot :D